OpenStudy (shamil98):

@Luigi0210 double integrals .-.

OpenStudy (shamil98):

\[S_n = \sum_{k=1}^{n} \] \[f(x_k,y_k) \Delta A_k \]

OpenStudy (shamil98):

If f is continous throughout R, then ,as you refine the mesh or two dimensional partition width to make both delta x and delta y go to zero, the sums in the first equation approach the limit called the double integral of f over R.

OpenStudy (shamil98):

\[\int\limits_{}^{} \int\limits_{R}^{} f(x,y) dA\]

OpenStudy (shamil98):

you could also write it as \[\int\limits_{}^{} \int\limits_{}^{R} f(x,y) dx~dy\]

OpenStudy (shamil98):

the R is on the bottom* i messed up with latex ..

OpenStudy (shamil98):

\[\int\limits_{}^{}\int\limits_{R}^{} f(x,y)~dA = \lim_{\Delta A \rightarrow 0}\sum_{k=1}^{n} f(x_k,y_k) \Delta A_k\]

OpenStudy (luigi0210):

This makes me miss calculus more ;-;

OpenStudy (shamil98):

what are you taking ? o.o

OpenStudy (luigi0210):

Nothing at all, which is why college is boring for me \[\large no math/science=boring=\not doing anything=\not doing hw=failing college=failing life=horrible life=hobo=death \]

OpenStudy (luigi0210):

*no math/science=boring=not doing anything=not doing hw=failing college=failing life=horrible life=hobo=death

OpenStudy (shamil98):

D:

OpenStudy (luigi0210):

True story bro

OpenStudy (shamil98):

dat sucks man, it'll get better .-.